Need a help from 7 through 12
Answer #1

7) B induced will be out of the plane
so I will be ccw so d)

8) V = - L dI/dt

so dI/dt = -V/L = -2/0.2= - 10

so a)

9) all voltage will be across the inductor so 20 V so e)

10)

I = V/R ( 1-e^(-t/tau))

tau = L/R

I = 4*(1-e^(-5.0E-3/(5.0E-3))= 2.5 A so c)

11)

emf = N B A w = 6*0.2*0.04*2*pi*50= 15 V so d)

12) Z = sqrt( R^2 + (w L - 1/( w C))^2)

= sqrt(1000^2 + (500*1 - 1/(500*1.0E-6))^2)= 1803 ohms

Imax = V/Z = 140/1803

V inductor rms = I max w L/sqrt(2) = (140/1803)*500*1/sqrt(2)= 27.5 V

so b)
Answer #2

d)counterclockwise

c)it is increasing at 10 A/s

d)20V